# Scalable Formal Verification via Autoencoder Latent Space Abstraction

### v2

@@ A.2. Proof of Lemma 3 (connected inclusion dynamics)
where $L_{y}:=\{u\in\mathbb{R}^{t}:A_{2}(u)=y\}$ is an affine subspace of $\mathbb{R}^{t}$. Observe that $M$ is the continuous image of the affine space TSTS under a coordinate-wise strictly monotone map, hence $M$ is a convex set defined with intervals over the first $n$ dimensions and constants for the next tnt-n dimensions. The intersection of convex sets is convex and hence connected. Therefore $f$ has connected pre-images.
@@ A.3. Proof of Lemma 6 (Conservative Labeling)
Recall $Z$ defines an under-approximation of domain $X$ in the latent space. Also recall that Eq. (9b) defines a high confidence over-approximation in the latent space for region $r_{i}$ by Proposition 5, that is $\hat{r_{Z,i}}$ defines a convex set in the latent space that will contain all points $\{h_{enc}{x}\mid x\in r_{i}\}$ with high confidence, therefore $Z\setminus\hat{r_{Z,i}}$ is an under-approximation of $X\setminus r_{i}$ in the latent domain $Z$. Also recall that in the original space $p_{n,i}\in L(x\in X\setminus r_{i})$, then by labeling $Z\setminus\hat{r_{Z,i}}$ with $p_{n,i}$ we under-approximate the label in the latent space. The logic follows similarly for labels $p_{i}$, as we first over-approximate $X\setminus r_{i}$ in the latent space. This implies that if $p_{i}\in L(q_{x})$ then $p_{i}\in L(x)$ for $h_{enc}(x)\in q_{x}$, hence $L(q_{x})\subseteq L(x)$.
@@ A.5. Proof of Theorem 11 (Sound Abstraction)
Assume $\omega_{x_{0}}\not\models\overline{\varphi}$, i.e. $\omega_{x_{0}}\models\neg\overline{\varphi}$. We begin with the base case $\overline{\varphi}:=p$. Then $\omega_{x_{0}}\not\models p$ means $p\notin L(x_{0})$ and since $L(q_{i})\subseteq L(x_{i})$ this implies $p\notin L(q_{0})$ and therefore $\omega_{q_{0}}\models\neg p$. Hence satisfaction of negative propositions in APAP is preserved from System (1) to $\mathcal{N}$. Since $\omega_{q_{0}}$ remains in relation with $\omega_{x_{0}}$, the satisfaction of negated propositions holds at subsequent states. By the semantics of LTL this satisfaction is applied inductively to Boolean and Temporal Operators. Then $\omega_{x_{0}}\not\models\overline{\varphi}$ implies $\exists\omega_{q_{0}}\not\models\overline{\varphi}$. This establishes the contrapositive and hence $\Omega_{q}\models\overline{\varphi}\implies\omega_{x}\models\overline{\varphi}$.

### v3

@@ A.2. Proof of Lemma 3 (connected inclusion dynamics)
where $L_{y}:=\{u\in\mathbb{R}^{t}:A_{2}(u)=y\}$ is an affine subspace of $\mathbb{R}^{t}$. Observe that $M$ is the continuous image of the affine space $TS$ under a coordinate-wise strictly monotone map, hence $M$ is a convex set defined with intervals over the first $n$ dimensions and constants for the next tnt-n dimensions. The intersection of convex sets is convex and hence connected. Therefore $f$ has connected pre-images.
@@ A.3. Proof of Lemma 6 (Conservative Labeling)
Recall $Z$ defines an under-approximation of domain $X$ in the latent space. Also recall that Eq. (9b) defines a high confidence over-approximation in the latent space for region $r_{i}$ by Proposition 5, that is $\hat{r_{Z,i}}$ defines a convex set in the latent space that will contain all points $\{h_{enc}{x}\mid x\in r_{i}\}$ with high confidence, therefore $Z\setminus\hat{r_{Z,i}}$ is an under-approximation of $X\setminus r_{i}$ in the latent domain $Z$. Also recall that in the original space $p_{n,i}\in L(x\in X\setminus r_{i})$, then by labeling $Z\setminus\hat{r_{Z,i}}$ with $p_{n,i}$ we under-approximate the label in the latent space. The logic follows similarly for labels $p_{i}$, as we first over-approximate $X\setminus r_{i}$ in the latent space. This implies that if $p_{i}\in L(q_{x})$ then $p_{i}\in L(x)$ for $h_{enc}(x)\in q_{x}$, hence $L(q_{x})\subseteq L(x)$. Since each atomic proposition is under-approximated with confidence $\geq 1-\delta$ by over-approximating its negation then, by union bound, the label $L(q_{x})\subseteq L(x)$ holds with confidence $\geq(1-\delta)^{|AP|-|L(x)|}$.
@@ A.5. Proof of Theorem 11 (Sound Abstraction)
Assume $\omega_{x_{0}}\not\models\overline{\varphi}$, i.e. $\omega_{x_{0}}\models\neg\overline{\varphi}$. We begin with the base case $\overline{\varphi}:=p$. Then $\omega_{x_{0}}\not\models p$ means $p\notin L(x_{0})$ and since $L(q_{i})\subseteq L(x_{i})$ this implies $p\notin L(q_{0})$ and therefore $\omega_{q_{0}}\models\neg p$. Hence satisfaction of negative propositions in $AP$ is preserved from System (1) to $\mathcal{N}$. Since $\omega_{q_{0}}$ remains in relation with $\omega_{x_{0}}$, the satisfaction of negated propositions holds at subsequent states. By the semantics of LTL this satisfaction is applied inductively to Boolean and Temporal Operators. Then $\omega_{x_{0}}\not\models\overline{\varphi}$ implies $\exists\omega_{q_{0}}\not\models\overline{\varphi}$. This establishes the contrapositive and hence $\Omega_{q}\models\overline{\varphi}\implies\omega_{x}\models\overline{\varphi}$.